A dart hits the square dartboard shown below at a random point. Find the probability that the dart lands in the shaded circular region. Each side of the dartboard is 11 in , and the radius of the shaded region is 4 in.
Use the value 3.14 for π . Round your answer to the nearest hundredth.
1
Expert's answer
2020-12-01T06:22:15-0500

By definition, probability is a ratio of events from sample space that satisfy given condition to the whole sample space. So,

"\displaystyle P(A) = \frac{A_{circle}}{A_{square}} = \frac{\pi r^2}{s^2} = \frac{3.14 \times 16}{121} = 0.415"

Answer: 0.415